OpenStudy (anonymous):

HELP!! The second side of a triangular deck is 5 feet longer than the shortest side and a third side that is 5 feet shorter than twice the length of the shortest side. If the perimeter of the deck is 80 feet, what are the lengths of the three sides?

OpenStudy (anonymous):

and what are you having trouble with? creating the proper equations from the given statements?

OpenStudy (anonymous):

yes i dont know how to start?

OpenStudy (anonymous):

start with defining your variables

OpenStudy (anonymous):

what variables will you be using?

OpenStudy (anonymous):

x i guess should it be x, 5-2x, and 5+x?

OpenStudy (anonymous):

what is x? defining is basically stating what each variable will represent ex. let "x" be the length of the shortest side of the triangle let "x2" be the length of the 2nd side of the triangle and let "x3" be the length of the 3rd side of the triangle

OpenStudy (anonymous):

the statements you have written are "almost" correct (one of them is incorrect) but they have absolutely no meaning because you did not state what is what

OpenStudy (anonymous):

Anyways starting with the first statement The second side of a triangular deck is 5 feet longer than the shortest side second side "x2" is "=" 5 feet longer than the shortest side "x+5" thus giving you the equation x2 = x+5

OpenStudy (anonymous):

moving onto the 2nd statement third side that is 5 feet shorter than twice the length of the shortest side. third side "x3" that is "=" 5 feet shorter than twice the length of the shortest side 5 feet smaller than twice the smallest side will be written as "2x-5" the resulting equation would be x3= 2x-5

OpenStudy (anonymous):

any questions at this point?

OpenStudy (anonymous):

um no thank you very much^^ I get it now

OpenStudy (anonymous):

would you like me to finish or you can solve the rest on your own?

OpenStudy (anonymous):

um would you help me finish please lol

OpenStudy (anonymous):

ok now, you have x - length of shortest side x2= x+5 and x3 = 2x-5 now, we're told that the perimeter of the triangle is 80 feet now, perimeter means?

OpenStudy (anonymous):

the total of all sides

OpenStudy (anonymous):

right, which also means we can write it as the equation P= x + x2 + x3 where P= 80 x2 = x+5 and x3= 2x-5 now, just substitute all the given information and you'll get 80 = x +(x+5) +(2x-5)

OpenStudy (anonymous):

and solve for x

OpenStudy (anonymous):

ohhh ok so i would solve for x and substitute it into the other equations

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok thank you very much for a very detailed explanation! ;_;

OpenStudy (anonymous):

no problem